### v1

# Reconstruction of Caterpillar Tanglegrams

Abstract.
A tanglegram consists of two rooted binary trees with the same number of leaves and a perfect matching between the leaves of the trees. Given a size-$n$ tanglegram, i.e., a tanglegram for two trees with $n$ leaves, a multiset of induced size-$(n-1)$ tanglegrams is obtained by deleting a pair of matched leaves in every possible way. Here, we analyze whether a size-$n$ tanglegram is uniquely encoded by this multiset of size-$(n-1)$ tanglegrams. We answer this question affirmatively in the case that at least one of the two trees of the tanglegram is a caterpillar tree.
Keywords: tanglegram, rooted binary tree, caterpillar, graph reconstruction
2020 Mathematics Subject Classification: 05C05, 05C10, 05C60
1. Introduction
A tanglegram $\mathcal{T}$ consists of two rooted binary trees $L$ and $R$ with the same number of leaves and a perfect matching $\sigma$ between the two leaf sets. In biology, tanglegrams are used in the study of cospeciation and coevolution. For example, tree $L$ may represent the phylogeny of a host, tree $R$ the phylogeny of a parasite, and the matching $\sigma$ the connection between the host and the parasite [11, 17]. Tanglegrams lead to a variety of mathematical questions, including their enumeration [1, 6, 13], drawing them with the fewest number of crossings possible [3, 5, 17], and their typical shape [9].
In this paper, we study the tanglegram reconstruction problem originally posed by Stephan Wagner and László Székely in 2017: Given a tanglegram $\mathcal{T}$ of size $n$ (where $n$ is the number of leaves in $L$ or $R$), does the multiset $\mathcal{D}^{m}(\mathcal{T})$ of induced size-$(n-1)$ tanglegrams obtained from $\mathcal{T}$ by deleting a pair of matched leaves from $\mathcal{T}$ in every possible way uniquely determine $\mathcal{T}$ for $n$ sufficiently large? This problem is motivated by the famous graph reconstruction conjecture that states that every graph $G$ on at least three vertices is uniquely determined (up to isomorphism) from the multiset of graphs obtained by deleting one vertex in every possible way from $G$ [8, 16]. This conjecture has been proven for many special classes of graphs but remains open in general (see [2, 7, 14] for an overview). It has also motivated various other reconstruction problems such as for partially ordered sets [10], permutations [15], and groups [12].
Here, we affirmatively answer the tanglegram reconstruction problem for the case that at least one of the two trees of the tanglegram is a caterpillar tree, where a caterpillar tree is a tree whose internal vertices form a path, such that the root of the tree is an endpoint of the path. We establish the following main theorem.
Theorem 1.1.
Let $\mathcal{T}=(L,R,\sigma)$ be a tanglegram of size $n\geq 5$. If $L$ or $R$ is a caterpillar, then $T$ is reconstructable from the multideck $\mathcal{D}^{m}(\mathcal{T})$.
The remainder of the paper is organized as follows. In Section 2, we introduce general terminology and notation. We then define different types of trees in Section 3. In Section 4, we consider the tanglegram reconstruction problem involving caterpillar trees and prove Theorem 1.1.
2. Preliminaries
We begin by introducing terminology and notation.
Definition 2.1.
A rooted tree is a tree $T$ where a special vertex, denoted $r_{T}$, is identified as the root of $T$. The parent of a non-root vertex $v$ is $v$’s neighbor on the $r_{T}-v$ path in $T$ and $y$ is a child of $v$ if $v$ is the parent of $y$. The leaves of $T$ are the vertices with no children. Two leaves with a common parent are called a cherry. We let $\mathbb{L}(T)$ denote the set of leaves and $|T|$ denote the number of leaves. We refer to $|T|$ as the size of $T$.
Definition 2.2.
A rooted binary tree is a rooted tree in which every vertex has zero or two children.
We often simply refer to this as a tree, as we will not consider other types of trees. Note that in a rooted binary tree we have that either the root is a leaf (and the tree has size 1) or the root has degree 2 and non-root internal vertices have degree 3.
We remark that whenever we write $T=T^{\prime}$, we mean that $T$ and $T^{\prime}$ are equal up to an isomorphism that maps root to root. In general, a rooted binary tree can be drawn in the plane in multiple ways, as shown in Figure 1
Figure 1. Two drawings for the same rooted binary tree $T$.
Definition 2.3.
Let $T_{1},T_{2}$ be rooted binary trees and let $T_{1}\oplus T_{2}$ be the rooted binary tree $T$ obtained by taking a new vertex $r_{T}$ and joining the roots of $T_{1}$ and $T_{2}$ to $r_{T}$. If $T=T_{1}\oplus T_{2}$, we say $T$ is composed of $T_{1}$ and $T_{2}$. Furthermore, we call $T_{1}$ and $T_{2}$ the maximal pending subtrees of $T$.
Since we consider unlabeled rooted binary trees up to isomorphism, the operation $\oplus$ is commutative. However, if not stated otherwise, for $T=T_{1}\oplus T_{2}$, we assume $|T_{1}|\leq|T_{2}|$.
Definition 2.4.
The caterpillar on $n$ leaves, denoted $C_{n}$, is the rooted binary tree $C_{n}=C_{1}\oplus C_{n-1}$, where $C_{1}$ is the unique rooted binary tree of size 1.
Alternatively, one can view the caterpillar of size $n$ as the unique tree in which the removal of all leaves results in the path graph on $n-1$ vertices, such that the root of the tree is an endvertex of the path. Note that the tree in Figure 1 is $C_{5}$.
We now turn to induced subtrees and decks.
Definition 2.5.
Let $T$ be a rooted binary tree with leaf set $\mathbb{L}(T)$ and let $S\subseteq\mathbb{L}(T)$. The rooted binary tree induced by $S$, denoted $T[S]$, is the tree formed by starting with the minimal subtree of $T$ containing all the vertices in $S$, declaring as root its closest vertex to the root of $T$, and then suppressing all internal vertices with only one child. We also refer to $T[S]$ as an induced subtree of $T$.
Definition 2.6.
The deck $\mathcal{D}(T)$ of $T$ is the set of trees $T^{\prime}$ with $|T^{\prime}|=|T|-1$ such that $T^{\prime}$ is an induced subtree of $T$. The multideck $\mathcal{D}^{m}(T)$ of $T$ is the multiset where the multiplicity of every $T^{\prime}\in\mathcal{D}(T)$ is |$\left|\{S\subseteq\mathbb{L}(T):T^{\prime}=T[S]\}\right|$.
We often refer to the elements of $\mathcal{D}^{m}(T)$ (respectively $\mathcal{D}(T)$) as cards. We say that a size-$n$ tree $T$ is reconstructable from its deck (resp. multideck) if whenever $T_{1}$ is a size-$n$ tree with $T_{1}\neq T$, we have $\mathcal{D}(T)\neq\mathcal{D}(T_{1})$ (resp. $\mathcal{D}^{m}(T)\neq\mathcal{D}^{m}(T_{1})$).
The following result shows that for sufficiently large trees, trees are reconstructable from decks and multidecks.
Theorem 2.7 ([4]).
For any tree of size $n\geq 5$, $T$ is reconstructable from the multideck $\mathcal{D}^{m}(T)$. If $n\geq 6$, then $T$ is reconstructable from the deck $\mathcal{D}(T)$.
We are now in the position to define the central concept of this paper.
Definition 2.8.
A tanglegram $\mathcal{T}=(L,R,\sigma)$ consists of two rooted binary trees, the left tree $L$ and the right tree $R$, along with a perfect matching $\sigma$ between their leaves. We will denote elements of $\sigma$ using edges $vu$, where $v\in\mathbb{L}(L)$ and $u\in\mathbb{L}(R)$. For an edge $vu\in\sigma$, we also say that $v$ and $u$ are matched. The size of a tanglegram is the common size of its underlying trees.
An isomorphism between two tanglegrams is a graph isomorphism that maps the root of the left (right) tree to the root of the left (right) tree. A tanglegram layout is a straight line drawing in which the two trees are drawn without crossing on the two sides of a strip and the matching edges are drawn inside the strip. Tanglegrams can have several layouts, as shown in Figure 2.
Figure 2. Two drawings of the same tanglegram.
We next define induced subtanglegrams of a tanglegram $\mathcal{T}=(L,R,\sigma)$, as well as the (multi)deck of a tanglegram.
Definition 2.9.
Let $\mathcal{T}=(L,R,\sigma)$ be a tanglegram. For any $E\subseteq\sigma$, let $S_{L}\subseteq\mathbb{L}(L)$ and $S_{R}\subseteq\mathbb{L}(R)$ be the set of leaves matched by $E$. The subtanglegram $\mathcal{T}[E]$ induced by $E$ is defined as the tanglegram formed from the induced subtrees $L[S_{L}]$ and $R[S_{R}]$ with the matching given by $E$.
When $E=\sigma-\{uv\}$ for some $vu\in\sigma$, we abuse notation and also denote $\mathcal{T}[\sigma-\{uv\}]$ as $\mathcal{T}-vu$. If the focus is on a particular leaf removed from $L$ (respectively $R$), then we abuse notation and also denote this as as $\mathcal{T}-v$ (respectively $\mathcal{T}-u$).
Definition 2.10.
The deck $\mathcal{D}(\mathcal{T})$ of the tanglegram $\mathcal{T}$ is the set of size $|\mathcal{T}|-1$ tanglegrams that are induced subtanglegrams of $\mathcal{T}$. The multideck $\mathcal{D}^{m}(\mathcal{T})$ of the tanglegram $\mathcal{T}$ contains each element $\mathcal{T}^{\prime}\in\mathcal{D}(\mathcal{T})$ with multiplicity |$|\{E\subseteq\sigma:\mathcal{T}^{\prime}=\mathcal{T}[E]\}|$.
As for trees, we often refer to the elements of $\mathcal{D}^{m}(\mathcal{T})$ (respectively $\mathcal{D}(\mathcal{T})$) as cards. An example is shown in Figure 3.
$\mathcal{T}$
--$\mathcal{T}_{1}$--$\mathcal{T}_{2}$--$\mathcal{T}_{3}$
Figure 3. The deck of $\mathcal{T}$ is $\{\mathcal{T}_{1},\mathcal{T}_{2},\mathcal{T}_{3}\}$. The multideck of $\mathcal{T}$ contains two copies of $\mathcal{T}_{1}$, one copy of $\mathcal{T}_{2}$, and two copies of $\mathcal{T}_{3}$.
3. Types of trees
In this section, we partition the set of rooted binary trees into three types. This partition will be used in the next section to prove our main theorem.
Definition 3.1.
Let $i\in\mathbb{N}$ and $T_{1},T_{2}$ be rooted binary trees. We define the operations $(T_{1}\oplus)^{i}T_{2}$ as follows: $(T_{1}\oplus)^{0}T_{2}=T_{2}$ and for $i>0$, $(T_{1}\oplus)^{i}(T_{2})=T_{1}\oplus((T_{1}\oplus)^{i-1}T_{2})$.
Definition 3.2.
Let $T$ be a rooted binary tree. Then $T$ is strippable if $|T|\geq 2$ and $T=C_{1}\oplus T^{\prime}$ for some tree $T^{\prime}$. Otherwise, $T$ is non-strippable. The stripping of $T$ is $(i,Q)$, where $Q$ is a non-strippable tree, $i$ is a nonnegative integer, and $T=(C_{1}\oplus)^{i}Q$. We call the leaves from the $C_{1}$’s strippable in $T$. Note that $i=0$ indicates that $T=T^{\prime}$, and that $T$ is non-strippable.
Note that $C_{n}$ has stripping $(n-1,C_{1})$. When $T$ is strippable, we will often label the strippable leaves according to their distance from the root (the only time this is ambiguous is the two leaves of $C_{n}$ at distance $n-1$ from the root).
Definition 3.3.
Let $T$ be a strippable tree with stripping $(i,Q)$. A labeling of the strippable leaves with subscripted symbols $\{v_{1},\ldots,v_{i}\}$ is a distance-to-root labeling if $v_{k}$ is distance $k$ from $r_{T}$ for $1\leq k\leq i$. When $T=C_{n}$, the distance-to-root labeling gives the remaining vertex label $v_{n}$.
Note that when $T=C_{n}$, the labels $v_{n-1}$ and $v_{n}$ may be used in two ways (interchanging the labels between the two leaves involved) in a distance-to-root labeling. Throughout, we will often identify a leaf with its label. For example, if we say that $v_{i}$ is matched to some leaf $\ell$, we mean that the leaf labeled by $v_{i}$ is matched to $\ell$. Furthermore, we often simply refer to the label set $\{v_{1},\ldots,v_{i}\}$ as the distance-to-root labeling.
Definition 3.4.
Let $T$ be a rooted binary tree. We say $T$ is type 00 if it is a caterpillar, type $1$ if it is strippable and not a caterpillar, and type $2$ otherwis
Notice that if $T$ is type $1$ with stripping $(i,Q)$, then $|Q|\geq 4$ and $|T|=|Q|+i\geq 5$. If $T$ is type $2$, then $T=R_{1}\oplus R_{2}$ for some rooted binary trees $R_{1},R_{2}$ with $|R_{1}|,|R_{2}|\geq 2$.
In Figure 4, we depict the three rooted binary trees of size 5. They are of types 0, 1, and 2, respectively.
Figure 4. The three trees of size $5$, respectively of types 0, 1, and 2.
4. Reconstruction of caterpillar tanglegrams
In this section, we prove reconstruction for tanglegrams where at least one tree is a caterpillar.
Throughout, we freely use the fact that if $\mathcal{T}=(L,R,\sigma)$ is a tanglegram of size $n\geq 5$, the multideck $\mathcal{D}^{m}(\mathcal{T})$ uniquely determines the trees $L$ and $R$. This is a direct consequence of Theorem 2.7 and is formalized by the following corollary.
Corollary 4.1.
For any tanglegram $\mathcal{T}=(L,R,\sigma)$ of size $n\geq 5$, the trees $L$ and $R$ are reconstructable from the multideck $\mathcal{D}^{m}(\mathcal{T})$.
We start with the case when both trees are caterpillars and use the distance-to-root labelings $\{v_{1},\ldots,v_{n}\}$ and $\{u_{1},\ldots,u_{n}\}$ of $L$ and $R$, respectively. Furthermore, for a card $(L^{\prime},R^{\prime},\sigma^{\prime})\in\mathcal{D}^{m}(\mathcal{T})$, we use subscripted symbols $\{v_{1}^{\prime},\ldots,v_{n-1}^{\prime}\}$ (resp. $\{u_{1}^{\prime},\ldots,u_{n-1}^{\prime}\}$), for the distance-to-root labeling of the induced subtree $L^{\prime}$ (resp. $R^{\prime}$).
We first handle the case in which $v_{1}$ is matched to the cherry of $R$.
Lemma 4.2.
Let $\mathcal{T}=(C_{n},C_{n},\sigma)$ be a tanglegram of size $n\geq 5$, where the left tree has distance-to-root labeling $\{v_{1},\ldots,v_{n}\}$. If $\sigma$ matches $v_{1}$ with a leaf of the cherry of the right tree, then $\mathcal{T}$ is reconstructable from the multideck $\mathcal{D}^{m}(\mathcal{T})$.
Proof.
The case $n=5$ can be checked directly, so throughout, we assume $n\geq 6$.
We first prove that $\mathcal{D}^{m}(\mathcal{T})$ determines whether or not $v_{1}$ is matched to a leaf of the cherry of the right tree. Let $\{u_{1},\ldots,u_{n}\}$ be the distance-to-root labeling of the right tree. If $v_{1}$ is adjacent to a leaf of the cherry, we may assume without loss of generality that the distance-to-root labeling is such that $v_{1}$ is adjacent to $u_{n}$. Observe that if $v_{1}$ is matched to $u_{n}$, then at least $n-1$ cards in $\mathcal{D}^{m}(\mathcal{T})$ have $v_{1}^{\prime}$ matched to a leaf in the cherry of $R^{\prime}$. If $v_{1}$ is matched to $u_{n-2}$, then between two and three cards (corresponding to the removal of $u_{n-1},u_{n}$, and possibly $v_{1}$) have $v_{1}^{\prime}$ matched to the cherry of $R^{\prime}$. If $v_{1}$ is matched to some $u_{i}$ with $i\leq n-3$, then at most one card has $v_{1}^{\prime}$ matched to the cherry of $R^{\prime}$. Hence, by counting cards in which $v_{1}^{\prime}$ is matched to the cherry of $R^{\prime}$, we can determine from $\mathcal{D}^{m}(\mathcal{T})$ if $v_{1}$ is matched to $u_{n}$.
We now show that when $v_{1}$ is matched to $u_{n}$, $\mathcal{T}$ is reconstructable from $\mathcal{D}^{m}(\mathcal{T})$. There are various cases and subcases based on which vertices $v_{2}$ and potentially $v_{3}$ are matched to in $\mathcal{T}$. These are collected in Table 1. Observe that we can distinguish the case where $v_{2}$ is matched to $u_{j}$ with $j<n-2$, as there exists a unique card, $\mathcal{T}-v_{1}u_{n}$, where $v_{1}^{\prime}$ is not matched to the cherry of $R^{\prime}$. Using this card also allows us to reconstruct $\mathcal{T}$.
Table 1. Case analysis for $v_{1}u_{n}\in\sigma$. We assume without loss of generality that in the card $\mathcal{T}^{\prime}=(L^{\prime},R^{\prime},\sigma^{\prime})$, the distance-to-root labeling of $R^{\prime}$ is chosen so that $v_{1}^{\prime}$ is matched to $u_{n-1}^{\prime}$ whenever possible.
Case Subcase Card $\mathcal{T}^{\prime}$ Feature
$v_{2}u_{n-1}\in\sigma$ $v_{3}u_{n-2}\in\sigma$ all $v_{1}^{\prime}u_{n-1}^{\prime},v_{2}^{\prime}u_{n-2}^{\prime}\in\sigma^{\prime}$
$v_{3}u_{n-2}\notin\sigma$ $\mathcal{T}-v_{i},i\leq 2$ $v_{1}^{\prime}u_{n-1}^{\prime}\in\sigma^{\prime},\,v_{2}^{\prime}u_{n-2}^{\prime}\notin\sigma^{\prime}$
all remaining $v_{1}^{\prime}u_{n-1}^{\prime},v_{2}^{\prime}u_{n-2}^{\prime}\in\sigma^{\prime}$
$v_{2}u_{n-2}\in\sigma$ $v_{3}u_{n-1}\in\sigma$ $\mathcal{T}-v_{i},i\leq 3$ $v_{1}^{\prime}u_{n-1}^{\prime},v_{2}^{\prime}u_{n-2}^{\prime}\in\sigma^{\prime}$
all remaining $v_{1}^{\prime}u_{n-1}^{\prime}\in\sigma^{\prime},\,v_{2}^{\prime}u_{n-2}^{\prime}\notin\sigma^{\prime}$
$v_{3}u_{n-1}\notin\sigma$ $\mathcal{T}-u_{n-1}$ $v_{1}^{\prime}u_{n-1}^{\prime},v_{2}^{\prime}u_{n-2}^{\prime}\in\sigma^{\prime}$
all remaining $v_{1}^{\prime}u_{n}^{\prime}\in\sigma^{\prime},v_{2}^{\prime}u_{n-2}^{\prime}\notin\sigma^{\prime}$
all remaining None $\mathcal{T}-v_{1}$ $v_{1}^{\prime}u_{n-1}^{\prime}\notin\sigma^{\prime}$
all remaining $v_{1}^{\prime}u_{n-1}^{\prime}\in\sigma^{\prime}$
For the remaining four subcases, observe that the number of cards where $v_{1}^{\prime}$ and $v_{2}^{\prime}$ are matched to the cherry of $R^{\prime}$ is $n$ (if $v_{2}$ is matched to $u_{n-1}$ and $v_{3}$ is matched to $u_{n-2}$), $n-2$ (if $v_{2}$ is matched to $u_{n-1}$ but $v_{3}$ is not matched to $u_{n-2}$), $3$ (if $v_{2}$ is matched to $u_{n-2}$ and $v_{3}$ is matched to $u_{n-1}$), or $1$ (if $v_{2}$ is matched to $u_{n-2}$ but $v_{3}$ is not matched to $u_{n-1}$). As $n\geq 6$, these are all distinct, so $\mathcal{D}^{m}(\mathcal{T})$ allows us to distinguish these four subcases. It now suffices to show that we can reconstruct $\mathcal{T}$ in each of them.
In the case where $v_{2}u_{n-1},v_{3}u_{n-2}\in\sigma$, there is a maximal sequence $v_{1},\ldots,v_{m}$ of vertices satisfying $v_{i}u_{n-i+1}\in\sigma$ for all $1\leq i\leq m$, where $m\geq 3$ (see Figure 5).
[…]
Figure 5. The tanglegram $\mathcal{T}$ in case $v_{i}$ is matched to $u_{n-i+1}$ ($1\leq i\leq m$) and its cards $\mathcal{T}-v_{i}$, for $1\leq i\leq m$.
If $m=n$, then all cards are identical. If $m<n$, then the $m$ cards $\mathcal{T}-v_{i}u_{n-i+1}$ with $1\leq i\leq m$ are identical with $v_{1}^{\prime},\ldots,v_{m-1}^{\prime}$ satisfying $v_{i}^{\prime}$ being matched to $u_{n-i}^{\prime}$, while $v_{m}^{\prime}$ is not matched to $u_{n-m}^{\prime}$. The other $n-m$ cards in $\mathcal{D}^{m}(\mathcal{T})$ have $v_{i}^{\prime}$ matched to $u_{n-i^{\prime}}$ for $1\leq i\leq m$. Thus, we can determine $m$, and using one of the identical cards $\mathcal{T}-v_{i}u_{n-i+1}$ with $1\leq i\leq m$, we can reconstruct $\mathcal{T}$.
In the case where $v_{2}u_{n-1}\in\sigma$ and $v_{3}u_{n-2}\notin\sigma$, the two cards where $v_{2}^{\prime}$ is not matched to a leaf in the cherry of $R^{\prime}$ are $\mathcal{T}-v_{1}$ and $\mathcal{T}-v_{2}$. Furthermore, these two cards are identical, and either one of them can be used to reconstruct $\mathcal{T}$.
In the case where $v_{2}u_{n-2},v_{3}u_{n-1}\in\sigma$, the three cards where $v_{1}^{\prime}$ and $v_{2}^{\prime}$ are matched to the cherry of $R^{\prime}$ are all the same. These correspond to $\mathcal{T}-v_{1},\mathcal{T}-v_{2}$, and $\mathcal{T}-v_{3}$. Using any of these three cards allows us to reconstruct $\mathcal{T}$.
Finally, in the case where $v_{2}u_{n-2}\in\sigma$ and $v_{3}u_{n-1}\notin\sigma$, we have that $u_{n-1}$ is matched to some $v_{k}$ with $k>3$. There is a unique card, $\mathcal{T}-v_{k}u_{n-1}$, in which the cherry of $R^{\prime}$ is matched to both $v_{1}^{\prime}$ and $v_{2}^{\prime}$. Among the remaining cards, there are $n-k$ cards (corresponding to the removal of $v_{i}$ with $i>k$) where $v_{k}^{\prime}$ is matched to the cherry of $R^{\prime}$ and $k-1$ cards (corresponding to the removal of $v_{i}$ with $i<k$) where $v_{k-1}^{\prime}$ is matched to the cherry of $R^{\prime}$. Hence, we can identify $k$, and using the card $\mathcal{T}-v_{k}u_{n-1}$ allows us to reconstruct $\mathcal{T}$. ∎
Lemma 4.3.
Let $\mathcal{T}=(C_{n},C_{n},\sigma)$ be a tanglegram of size $n\geq 5$. Then $\mathcal{T}$ is reconstructable from $\mathcal{D}^{m}(\mathcal{T})$.
Proof.
Let $\{v_{1},\ldots,v_{n}\}$ be the distance-to-root labeling of the left tree $L=C_{n}$ and let $\{u_{1},\ldots,u_{n}\}$ be the distance-to-root labeling of the right tree $R=C_{n}$. Furthermore, for a card $(L^{\prime},R^{\prime},\sigma^{\prime})\in\mathcal{D}^{m}(\mathcal{T})$, let $\{v_{1}^{\prime},\ldots,v_{n-1}^{\prime}\}$ and $\{u_{1}^{\prime},\ldots,u_{n-1}^{\prime}\}$ denote the distance-to-root labelings of the induced subtrees $L^{\prime}$ and $R^{\prime}$.
Assume $v_{1}$ is matched to $u_{i}$ and $v_{2}$ is matched to $u_{j}$ with $i\neq j$. By Lemma 4.2, $\mathcal{T}$ is reconstructable if $i\geq n-1$, so we will assume $i<n-1$. Additionally, if $j\in\{n-1,n\}$, we set $j=n-1$.
If $i=1$, then at least $n-1$ of the cards in $\mathcal{D}^{m}(\mathcal{T})$ have $v_{1}^{\prime}$ matched to $u_{1}^{\prime}$. If instead $i\neq 1$, then there are at most two such cards (corresponding to $j=1$ or $i=2$). Thus, from $\mathcal{D}^{m}(\mathcal{T})$, we can determine whether or not $i=1$.
Suppose $i=1$. Let $v_{1},\ldots,v_{m}$ be a maximal sequence of vertices satisfying that $v_{i}$ is matched to $u_{i}$ for all $1\leq i\leq m$. If $m=n$, all cards are identical and the tanglegram is reconstructable. Since $m=n-1$ is not possible, we consider $m<n-1$. Then, in exactly $m$ identical cards (corresponding to $\mathcal{T}-v_{i}u_{i}$ with $1\leq i\leq m$), we have $v_{k}^{\prime}$ matched to $u_{k}^{\prime}$ for $11\leq k\leq m-$ 1 and $v_{m}^{\prime}$ not matched to $u_{m}^{\prime}$. In the remaining cards, $v_{k}^{\prime}$ and $u_{k}^{\prime}$ are matched for $1\leq k\leq m$. Thus, we can identify $m$, and the $m$ identical cards resulting from $\mathcal{T}-v_{i}u_{i}$ with $1\leq i\leq m$ allow us to reconstruct $\mathcal{T}$.
Suppose instead $i\neq 1$. In the cards of $\mathcal{D}^{m}(\mathcal{T})$, we consider which vertex is matched to $v_{1}^{\prime}$. From the cards other than $\mathcal{T}-v_{1}u_{i}$, we get $i-1$ cards with $v_{1}^{\prime}$ matched to $u_{i-1}^{\prime}$ (corresponding to the removal of $u_{k}$ with $1\leq k<i$) and $n-i$ cards with $v_{1}^{\prime}$ matched to $u_{i}^{\prime}$ (corresponding to the removal of $u_{k}$ with $i<k\leq n$). Which vertex $v_{1}^{\prime}$ is matched to on the remaining card, $\mathcal{T}-v_{1}u_{i}$, depends on the relationship of $i$ and $j$. The vertices matched to $v_{1}^{\prime}$ and their multiplicities in $\mathcal{D}^{m}(\mathcal{T})$ are collected in Table 2.
Table 2. For $21<i<n-$ 2, the vertices matched to $v_{1}^{\prime}$ in $\mathcal{D}^{m}(T)$ with multiplicities.
$j>i+1$ $j=i+1$ $j=i-1$ $j<i-1$
$u_{j-1}^{\prime}\times(1)$ $u_{i}^{\prime}\times(n-i+1)$ $u_{i}^{\prime}\times(n-i)$ $u_{i}^{\prime}\times(n-i)$
$u_{i}^{\prime}\times(n-i)$ $u_{i-1}^{\prime}\times(i-1)$ $u_{i-1}^{\prime}\times(i)$ $u_{i-1}^{\prime}\times(i-1)$
$u_{i-1}^{\prime}\times(i-1)$ $u_{j}^{\prime}\times 1$
From Table 2, if $v_{1}^{\prime}$ is matched to exactly two $u_{k}^{\prime}$, then $j=i+1$ or $j=i-1$ and the multiplicities allow us distinguish between these cases. If $j=i+1$, then there is a maximal sequence of vertices $v_{1},\ldots,v_{m}$ satisfying that $v_{k}$ is matched to $u_{i+k-1}$ for $11\leq k\leq m-$ 1. Hence, there are $m$ identical cards with $v_{k}^{\prime}$ matched to $u_{i+k-1}^{\prime}$ for $11\leq k\leq m-$ 1 and $v_{m}^{\prime}$ not matched to $u_{i+m-1}^{\prime}$. However, in all other cards, $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ are matched to a set of $u$’s that are consecutively and increasingly indexed. Thus, we can identify $m$, and one of the $m$ identical cards resulting from removing $v_{k}u_{i+k-1}$ with $1\leq k\leq m$ allows us to reconstruct $\mathcal{T}$. If instead $j=i-1$, then a similar argument with $v_{k}u_{i-k+1}\in\sigma$ for $1\leq k\leq m$ shows that $\mathcal{T}$ is reconstructable.
Finally, we consider when $v_{1}^{\prime}$ is matched to $u_{k}^{\prime}$ for three different values of $k$. If these indices are non-consecutive, then we can easily identify both $i$ and the card $\mathcal{T}-v_{1}u_{i}$. In this case, $\mathcal{T}$ is reconstructable from this card. If the indices are consecutive, then $j=i+2$ or $j=i-2$ and we consider the multiplicity of the cards in which $v_{1}^{\prime}$ is matched to $u_{k}^{\prime}$ with $k$ maximum. In the former case, there is one such card and in the latter case, there are $n-i\geq 2$ such cards (since $i\leq n-1$). Hence, we can distinguish between these cases, identify both $i$ and the card $\mathcal{T}-v_{1}u_{i}$, and reconstruct $\mathcal{T}$. ∎
We now consider the remaining cases where exactly one of $L$ or $R$ is a caterpillar. Without loss of generality, $L=C_{n}$. We first consider the case that $R$ is type 1.
Lemma 4.4.
Let $\mathcal{T}=(C_{n},R,\sigma)$ be a tanglegram of size $n\geq 5$ such that $R$ is type 1. Then, $\mathcal{T}$ is reconstructable from the multideck $\mathcal{D}^{m}(C_{n},R,\sigma)$.
Proof.
Let $\{v_{1},\ldots,v_{n}\}$ be a distance-to-root labeling of the leaves of $C_{n}$ and, on each card, let $\{v_{1}^{\prime},\ldots,v_{n-1}^{\prime}\}$ be a distance-to-root labeling of the leaves of $C_{n-1}$. Further, let $(i,Q)$ be the stripping of $R$ (with $|Q|\geq 4$ since $R$ is type 1). As Table 3 shows, $\mathcal{D}^{m}(\mathcal{T})$ allows us to determine whether both, none, or exactly one of $v_{n}$ and $v_{n-1}$ are matched to a leaf from $Q$.
Table 3. Case analysis for the number of cards in $\mathcal{D}^{m}(\mathcal{T})$ with the property that both of $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$, exactly one of them, or none of them is matched to a leaf of $Q$ depending on whether both of $v_{n}$ and $v_{n-1}$, exactly one of them, or none of them, is matched to a leaf of $Q$ in $\mathcal{T}$.
Case number of cards with both $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ matched to $Q$ number of cards with exactly one of $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ matched to $Q$ number of cards with none of $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ matched to $Q$
both $v_{n}$ and $v_{n-1}$ matched to $Q$ $\geq n-2\geq 3$ $\leq 2$
exactly one of $v_{n}$ and $v_{n-1}$ matched to $Q$ $\leq 1$ $\geq n-2\geq 3$ $\leq 1$
none of $v_{n}$ and $v_{n-1}$ matched to $Q$ $\leq 2$ $\geq n-2\geq 3$
We now consider these three cases individually:
  1. (1)
    First, assume that both $v_{n}$ and $v_{n-1}$ are matched to $Q$. Let $v_{n},v_{n-1},\ldots,v_{n-s}$ with $11\leq s\leq n-i-$ 1 be a maximal sequence of $s+1$ leaves all matched to $Q$, whereas $v_{n-s-1},\ldots,v_{n-t}$ with $s+1\leq t\leq n-1$ is a maximal sequence of $t-s\geq 1$ leaves not matched to $Q$.
    In this case, there are precisely $s+1$ cards (corresponding to the removal of $v_{n},v_{n-1},\ldots,v_{n-s}$, respectively) with the property that there is a maximal sequence of $s$ leaves $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s}$ all matched to $Q$, whereas $v^{\prime}_{n-s-1},\ldots,v^{\prime}_{n-t}$ is a maximal sequence of $t-s$ leaves not matched to $Q$ (on the remaining $n-s-1\geq 1$ cards, at least $s+1$ leaves $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s-1},\ldots$ are matched to $Q$). Thus, we can determine $s$ and $t$.
    Now, using a card containing a maximal sequence of $s$ leaves $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s}$ matched to $Q$ (meaning that a leaf from $Q$ was removed but all strippable leaves of $R$ are present), allows us to determine which of the leaves $v_{1},\ldots,v_{n-s-1}$ are matched to strippable leaves of $R$ and which are matched to $Q$. Furthermore, for those matched to strippable leaves of $R$, we can determine the matching.
    A second card with a maximal sequence of at least $s+1$ leaves
    $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s-1},\ldots$ all matched to $Q$ and a maximal sequence of $t-s-1\geq 0$ of the next (lower-indexed) leaves not matched to $Q$ (meaning that a strippable leaf was removed but all leaves from $Q$ are present), allows us to determine the matching for all remaining leaves. Thus, we can reconstruct $\mathcal{T}$ in this case.
  2. (2)
    Second, assume that none of $v_{n}$ and $v_{n-1}$ is matched to $Q$. In this case, $v_{n}$ and $v_{n-1}$ are both matched to strippable leaves of $R$. Thus, there is a maximal sequence of $s+1$ leaves $v_{n},v_{n-1},\ldots,v_{n-s}$ matched to strippable leaves of $R$ and a maximal sequence of $t-s\geq 1$ leaves $v_{n-s-1},\ldots,v_{n-t}$ matched to $Q$. We can thus repeat the argument used in Case (1) with the roles of $Q$ and the strippable leaves of $R$ interchanged.
  3. (3)
    Finally, assume that exactly one of $v_{n}$ and $v_{n-1}$ is matched to $Q$. Without loss of generality assume that $v_{n}$ is matched to $Q$ and $v_{n-1}$ is matched to a strippable leaf of $R$. Now notice that if $v_{n-2}$ is matched to $Q$, then there is precisely one card with $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ both matched to $Q$. If $v_{n-2}$ is matched to a strippable leaf of $R$, then there is no such card. Thus, we can distinguish these two subcases using $\mathcal{D}^{m}(\mathcal{T})$.
    1. (a)
      If $v_{n-2}$ is matched to $Q$, there is a unique card, $\mathcal{T}-v_{n-1}$, with the property that both of $v^{\prime}_{n-1}$ and $v^{\prime}_{n-2}$ are matched to $Q$. Furthermore, this card has the property that all leaves of $Q$ are present and we can use it to determine the matching of all leaves of $C_{n}$ that are matched to $Q$. Now, any other card with the property that $R^{\prime}$ has at least $i$ strippable leaves (meaning that a leaf from $Q$ was removed) allows us to determine the matching of all leaves of $C_{n}$ matched to the strippable leaves of $R$. Combining these cards, we can reconstruct $\mathcal{T}$.
    2. (b)
      If $v_{n-2}$ is matched to a strippable leaf of $R$, let $v_{n-1},v_{n-2},\ldots,v_{n-t}$ with $12\leq t\leq i+$ 1 be a maximal sequence of $t$ leaves all matched to strippable leaves of $R$, whereas $v_{n-t-1}$ is matched to $Q$. Using similar reasoning as in the previous cases, we can determine $t$.
      Now, since $|Q|\geq 4$, we can find a card such that $R^{\prime}$ has at least $i$ strippable leaves (implying that a leaf from $Q$ has been removed), $v^{\prime}_{n-1}$ is matched to $Q$, and there is a maximal sequence of at least $t$ leaves $v^{\prime}_{n-2},v^{\prime}_{n-3},\ldots,v^{\prime}_{n-t-1},\ldots$ all matched to strippable leaves. We can use this card to determine the indices (in the distance-to-root labeling) and matching of the leaves of $C_{n}$ matched to the strippable leaves of $R$. Furthermore, we can identify a second card with $v_{n-1}^{\prime}$ matched to $Q$ and a maximal sequence of only $t-1$ leaves $v_{n-2}^{\prime},\ldots,v^{\prime}_{n-t}$ matched to strippable leaves of $R$ (implying that a strippable leaf has been removed while all leaves of $Q$ are present). This card allows us to determine the matching for all remaining leaves and we can reconstruct $\mathcal{T}$.
We next consider the case that $R$ is type 2.
Lemma 4.5.
Let $\mathcal{T}=(C_{n},R,\sigma)$ be a tanglegram of size $n\geq 5$ such that $R$ is type 2. Then, $\mathcal{T}$ is reconstructable from the multideck $\mathcal{D}^{m}(C_{n},R,\sigma)$.
Proof.
Let $\{v_{1},\ldots,v_{n}\}$ be a distance-to-root labeling of the leaves of $C_{n}$ and, on each card, let $\{v_{1}^{\prime},\ldots,v_{n-1}^{\prime}\}$ be a distance-to-root labeling of the leaves of $C_{n-1}$. Let $R=R_{1}\oplus R_{2}$ with $|R_{1}|,|R_{2}|\geq 2$ (since $R$ is type 2). As $n\geq 5$, we can reconstruct $C_{n}$ and $R$ from $\mathcal{D}^{m}(\mathcal{T})$ (Theorem 2.7).
We first argue that we can determine from $\mathcal{D}^{m}(\mathcal{T})$ whether $v_{n}$ and $v_{n-1}$ are matched to the same maximal pending subtree $R_{i}$ of $R$ or not. If $v_{n}$ and $v_{n-1}$ are matched to the same maximal pending subtree of $R$, then on at least $n-2\geq 3$ cards, leaves $v^{\prime}_{n-2}$ and $v^{\prime}_{n-1}$ are matched to the same maximal pending subtree of $R^{\prime}$. If $v_{n}$ and $v_{n-1}$ are not matched to the same maximal pending subtree of $R$, then on precisely one card, leaves $v^{\prime}_{n-2}$ and $v^{\prime}_{n-1}$ are matched to the same maximal pending subtree of $R^{\prime}$. Thus, we can distinguish these two cases.
  1. (1)
    First, assume that $v_{n}$ and $v_{n-1}$ are matched to the same maximal pending subtree of $R$. Let $v_{n},v_{n-1},\ldots,v_{n-s}$ with $31\leq s\leq n-$ 3 be a maximal sequence of $s+1$ leaves all matched to the same maximal pending subtree of $R$, say $R_{i}$, whereas $v_{n-s-1},\ldots,v_{n-t}$ with $s+1\leq t\leq n-1$ is a maximal sequence of $t-s\geq 1$ leaves matched to $R_{3-i}$.
    In this case, there are precisely $s+1$ cards (corresponding to the removal of $v_{n},v_{n-1},\ldots,v_{n-s}$, respectively) with the property that there is a maximal sequence of $s$ leaves $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s}$ all matched to $R_{i}$, whereas $v^{\prime}_{n-s-1},\ldots,v^{\prime}_{n-t}$ is a maximal sequence of $t-s$ leaves matched to $R_{3-i}$ (on the remaining $n-s-1\geq 1$ cards, at least $s+1$ leaves $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s-1},\ldots$ are matched to $R_{i}$). Thus, we can determine $s$ and $t$.
    Now, using a card containing a maximal sequence of $s$ leaves $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s}$ matched to $R_{i}$ (meaning that a leaf from $R_{i}$ was removed but all leaves from $R_{3-i}$ are present), allows us to determine which of the leaves $v_{1},\ldots,v_{n-s-1}$ are matched to $R_{3-i}$ and which are matched to $R_{i}$. Furthermore, for those matched to $R_{3-i}$, we can determine the matching.
    A second card with a maximal sequence of at least $s+1$ leaves
    $v^{\prime}_{n-1},\ldots,v^{\prime}_{n-s-1},\ldots$ all matched to $R_{i}$ and a maximal sequence of $t-s-1$ of the next (lower-indexed) leaves matched to $R_{3-i}$ (meaning that a leaf from $R_{3-i}$ was removed but all leaves from $R_{i}$ are present), allows us to determine the matching for all remaining leaves. Thus, we can reconstruct $\mathcal{T}$ in this case.
  2. (2)
    Second, assume that $v_{n}$ and $v_{n-1}$ are matched to different maximal pending subtrees of $R$. Without loss of generality assume that $v_{n}$ is matched to $R_{i}$ and $v_{n-1},v_{n-2}$ are matched to $R_{3-i}$. In particular, let $v_{n-1},v_{n-2},\ldots,v_{n-t}$ with $22\leq t\leq n-$ 2 be a maximal sequence of $t$ leaves all matched to $R_{3-i}$, whereas $v_{n-t-1}$ is matched to $R_{i}$. Using similar reasoning as in the previous case, we can determine $t$.
    Now, there is a unique card, $\mathcal{T}-v_{n}$, with a maximal sequence of $t$ leaves $v^{\prime}_{n-1},v^{\prime}_{n-2},\ldots,v^{\prime}_{n-t}$ matched to $R_{3-i}$ and we can use it to determine the indices (in the distance-to-root labeling) and matching of the leaves of $C_{n}$ matched to leaves from $R_{3-i}$, except for $v_{n-1}$ and $v_{n-2}$. Let $r$ denote the size of the subtree this sequence of $t$ leaves is matched to on this card. To resolve the matching for $v_{n-1}$ and $v_{n-2}$, we identify a second card with the property that $v^{\prime}_{n-1}$ is matched to $R_{i}$, there is a maximal sequence of $t$ leaves $v^{\prime}_{n-2},v^{\prime}_{n-3},\ldots,v^{\prime}_{n-t-1}$ matched to $R_{3-i}$, and $|R_{3-i}|=r$ (implying that a leaf from $R_{i}$ not matched to $v_{n}$ was removed; such a card must exist since $|R_{i}|\geq 2$). This card allows us to determine the matching of $v_{n-1}$ and $v_{n-2}$. Finally, we can identify a third card with $v_{n-1}^{\prime}$ matched to $R_{i}$ and a maximal sequence of only $t-1$ leaves $v_{n-2}^{\prime},\ldots,v^{\prime}_{n-t}$ matched to $R_{3-i}$ (implying that a leaf from $R_{3-i}$ was removed). This card allows us to determine the matching for all remaining leaves and we can reconstruct $\mathcal{T}$.
Lemmas 4.3, 4.4, and 4.5 establish our main theorem, which we re-state for the reader’s convenience.
Theorem 1.1.
Let $\mathcal{T}=(L,R,\sigma)$ be a tanglegram of size $n\geq 5$. If $L$ or $R$ is a caterpillar, then $T$ is reconstructable from the multideck $\mathcal{D}^{m}(\mathcal{T})$.
We end this manuscript by noting that Theorem 1.1 establishes the reconstruction of caterpillar tanglegrams of size $n\geq 5$ from the multideck $\mathcal{D}^{m}(\mathcal{T})$. Indeed, our proofs rely on the multiplicities of cards in several instances. However, since trees of size $n\geq 6$ are reconstructable from their decks (Theorem 2.7), it is an interesting open question to investigate whether decks also determine sufficiently large tanglegrams (note that we must have $n\geq 6$ for the trees to be reconstructable from the deck).
Problem 4.6.
Determine whether sufficiently large tanglegrams are reconstructable from their decks.
Acknowledgements
This project was initiated when the authors participated in the Mathematics Research Community workshop “Trees in Many Contexts” (Beaver Hollow, NY), organized by the American Mathematical Society with the support of the National Science Foundation under Grant Number DMS-1641020. The work on this project was continued when EC, LS, and KW were in residence at the Institute for Computational and Experimental Research in Mathematics of Brown University (Providence, RI), during the “Theory, Methods, and Applications of Quantitative Phylogenomics” semester program supported by the National Science Foundation under Grant Number DMS-1929284.